# Cutting through buggy adversarial example defenses: fixing 1 line of code breaks Sabre

### v1

@@ Abstract
Sabre is a defense to adversarial examples that was accepted at IEEE S&P 2024. We first reveal significant flaws in the evaluation that point to clear signs of gradient masking. We then show the cause of this gradient masking: a bug in the original evaluation code. By fixing a single line of code in the original repository, we reduce Sabre’s robust accuracy to 0%. In response to this, the authors modify the defense and introduce a new defense component not described in the original paper. But this fix contains a second bug; modifying one more line of code reduces robust accuracy to below baseline levels.
@@ 1 Introduction
In this short paper we show that by changing a single line of code to fix a bug in the published evaluation code, the attack success rate increases to 100%. The authors responded to our attack by modifying the published defense code to introduce a new defense component that was not present in the original paper, thereby changing the defense to a different defense than was initially proposed. Unfortunately, this modification contains a second bug, and by modifying a second line of code we again completely break the defense. The authors claim to have had a fix to this second attack for six weeks, but we have been unable to verify these claims because the authors have not published code for this fix, nor have they described the necessary changes. When the authors provide this information we will update our paper with an analysis of any future modifications or corrections to the currently published defense implementation.
@@ 6 Our attack
We break Sabre twice.
@@ 6.2 A fix to our first break
Throughout the paper, Sabre claims to be “end-to-end differentiable to avoid gradient masking”, claims “innate differentiability of [their] defense mechanism”, and claims to “ensure end-to-end differentiability” [5]. But with this new changed proposed by the authors, this is no longer the case: discretization is the quinticential example of nondifferentiability. Indeed, gradient-based attacks are only effective when models are usefully differentiable. And yet, the paper states it verifies “the absence of gradient obfuscation”.
@@ 6.3 Our second break
On CIFAR-10, where the defense sets the number of decimals to 1 and so again discretizes each input pixel to $\{0,0.1,0.2,\dots,0.9,1.0\}$, we achieve an attack success rate of 100%. Unfortunately, this CIFAR-10 result is for a model we have trained ourselves (but using the code published by the authors), and not for a model trained by the authors themselves. When we informed the authors of this second attack, they stated that this fix was incomplete for CIFAR-10, and that the defense must be adjusted differently for this model. However the authors have been unable to provide a description of the necessary changes, an implementation of the code for the fix, or a pointer to where the necessary adjustments are described in the original paper.
@@ 8 Conclusion
Sabre is the second defense to be accepted at IEEE S&P in as many years that had clear signs of a flawed evaluation which should have been easily identified by qualified reviewers. As we have shown, these evaluation flaws are caused by one (or two) bugs in the original evaluation that can be fixed by modifying one (or two) lines of code.

### v2

@@ Abstract
Sabre is a defense to adversarial examples that was accepted at IEEE S&P 2024. We first reveal significant flaws in the evaluation that point to clear signs of gradient masking. We then show the cause of this gradient masking: a bug in the original evaluation code. By fixing a single line of code in the original repository, we reduce Sabre’s robust accuracy to 0%. In response to this, the authors modify the defense and introduce a new defense component not described in the original paper. But this fix contains a second bug; modifying one more line of code reduces robust accuracy to below baseline levels. After we released the first version of our paper online, the authors introduced another change to the defense; by commenting out one line of code during attack we reduce the robust accuracy to 0% again.
@@ 1 Introduction
In this short paper we show that by changing a single line of code to fix a bug in the published evaluation code, the attack success rate increases to 100%. The authors responded to our attack by modifying the published defense code to introduce a new defense component that was not present in the original paper, thereby changing the defense to a different defense than was initially proposed. Unfortunately, this modification contains a second bug, and by modifying a second line of code we again completely break the defense. We then released the first version of our paper online. After this initial release of this paper, the authors updated the code again with another fix; but this contained a third bug, and by commenting out a single line code when running the attack (and putting it back at test time), we reduced defense accuracy to 0%.
@@ 6 Our attack
We break Sabre three times.
@@ 6.2 A fix to our first break
Throughout the paper, Sabre claims to be “end-to-end differentiable to avoid gradient masking”, claims “innate differentiability of [their] defense mechanism”, and claims to “ensure end-to-end differentiability” [5]. But with this new changed proposed by the authors, this is no longer the case: discretization is the quintessential example of nondifferentiability. Indeed, gradient-based attacks are only effective when models are usefully differentiable. And yet, the paper states it verifies “the absence of gradient obfuscation”.
@@ 6.3 Our second break
On CIFAR-10, where the defense sets the number of decimals to 1 and so again discretizes each input pixel to $\{0,0.1,0.2,\dots,0.9,1.0\}$, we achieve an attack success rate of 100%. Unfortunately, this CIFAR-10 result is for a model we have trained ourselves (but using the code published by the authors), and not for a model trained by the authors themselves. When we informed the authors of this second attack, they stated that this fix was incomplete for CIFAR-10, and that the defense must be adjusted differently for this model. The following section will break this modified defense.
6.4 Our third break
The first version of this paper was released on May 6th 2024, and as described above, attacked a model we had trained ourselves because the authors were unable to provide a pretrained model. Our initial paper contained the disclaimer “When the authors provide this information we will update our paper with an analysis of any future modifications or corrections to the currently published defense implementation.” On May 24th and 27th, the authors made several corrections to the code, also released a pretrained CIFAR-10 model. On the same day (May 27th) we broke this third release of the code, and updated the arXiv version of this paper.
The authors’ 300-LoC diff makes many changes to many pieces of the code, again introducing changes not described in the original paper. We omit a more complete discussion surrounding these changes from this paper, but only remark that this is the second time the defense has been dramatically altered since it was initially released by the authors. We firmly believe that code for research papers should accurately reflect the experiments actually performed when writing the paper, and should not be altered in major ways after the fact. But for the purpose of our paper here it is not necessary to address this topic, because only one of the changes matters: the denoising forward pass has been updated to the following method:
1def forward(self, inputs, x0):
2 b, _, m, n = inputs.shape
3 c = self.in_channels
4
5 inputs = reshape_inputs(inputs)
6 x0 = reshape_inputs(x0)
7
8 inputs = torch.cat((inputs, x0), dim=1)
9
10 inputs = ((inputs*(v := ((1<<3)-1))).round() - ((inputs<0) & (inputs%1 != 0)).float())/v
11 outputs = self.layers(inputs)
12
13 if math.sqrt(m * n) > int(math.sqrt(m * n)):
14 outputs = outputs.reshape(b, c, -1)[:, :, :m * n].reshape(b, c, m, n)
15 return outputs
Note that line 10 of this code (again) completely discretizes the inputs—this time without any attempted smoothing. This causes the gradient to be identically zero, and so PGD makes no progress during attack. As a result, the defense appears to obtain 90% robust accuracy on CIFAR-10.
To break the defense, we simply comment out this line of code when running the attack, and then put the line back in when running the final evaluation. This is essentially equivalent to wrapping the code with a BPDA block, but even simpler: because the inputs are being discretized to one of eight different levels, it is sufficient to pretend (during the attack) that there is no discretization at all. Below is our complete break:
1diff --git a/models/denoise.py b/models/denoise.py
2index ba8154c..472fbec 100644
3--- a/models/denoise.py
4$+$ b/models/denoise.py
5@@ -38,7 +38,7 @@ class DenoisingCNN(nn.Module):
6 inputs = torch.cat((inputs, x0), dim=1)
7
8- inputs = ((inputs*(v := ((1<<3)-1))).round() - ((inputs<0) & (inputs%1 != 0)).float())/v
9+ #inputs = ((inputs*(v := ((1<<3)-1))).round() - ((inputs<0) & (inputs%1 != 0)).float())/v
10
11 outputs = self.layers(inputs)
With this single change, we reduce the defense accuracy on the paper’s official CIFAR-10 implementation with the provided CIFAR-10 model to $0\%$ robust accuracy. As can be seen from this attack, one must be exceptionally careful when introducing defense components that may cause gradient masking. Each of the three releases of this paper contained a different, but significant, gradient masking operation, despite the paper’s claims it “ensure[s] end-to-end differentiability”.
@@ 8 Conclusion
Sabre is the second defense to be accepted at IEEE S&P in as many years that had clear signs of a flawed evaluation which should have been easily identified by qualified reviewers. As we have shown, these evaluation flaws are caused by between one and three bugs in the original evaluation that can be fixed by modifying one to three lines of code.